Find the area bounded by the curve y^2-3x+3=0 and the line x=4.


1
Expert's answer
2021-03-30T16:02:29-0400

We can consider the sketch of the graph below



Substituting x=4 in the equation "y^2-3x+3=0"

We get "y^2-3(4)+3=0"

"y^2- 9=0"

"y^2=9"

"y=-3\space or +3"

Intersection points are A(4,-3) and B(4,3)

y is changing from -3 to 3

"Area=\int_a^b(Right \space curve-Left \space curve)dy"

"Area=\int_{-3}^3(4-(1+ \frac{y^2}{3}))dy"

"Area=\int_{-3}^3(4-1- \frac{y^2}{3})dy"

"Area=\int_{-3}^3(3- \frac{y^2}{3})dy"

"Area=\int_{-3}^33dy- \int_{-3}^3\frac{y^2}{3}dy"

"Area=3[y]_{-3}^3-\frac{1}{3}[\frac{y^3}{3}]_{-3}^3"

"Area=18-6=12" square units